A ferry boat on a river is sailing with a velocity of 55km/h going 45
1
Expert's answer
2021-10-01T13:27:25-0400

Suppose that


A ferry boat on a river is sailing with a velocity of 55 km/h going southeast with an angle of 45°. What are the horizontal and vertical components of the boat’s motion?


"v_h=v\cdot\cos45°=55\cdot\cos45°=38.89\ (m/s)"


"v_v=v\cdot\sin45°=55\cdot\sin45°=38.89\ (m/s)"